OpenStudy (anonymous):

There were 350 children enrolled in the Drama Acsdemy in 2009. 2/7 of the children were girls. Additional children joined the drama academy in 2010. For every 2 girls who joined the drama academy, 1boy joined the drama academy. If there was an equal number of boys & girls in 2010, how many boys joined the academy in that year ?

OpenStudy (anonymous):

no of children in 2009 = 350 No of girls in 2009 = 2/7*350 = 100 No of boys in 2009 = 250 Now, in 2010 No of boys = girls also 1 boy joined for every 2 girls so to put in equation, let boys joined be x in 2010 250+X = 100+2X solving gives, X = 150 No of boys joined is 150 in 2010.